OpenStudy (rebecca1588):

Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y A. The only solution is (24, 0). B. There is no solution. C. There are an infinite number of solutions. D. The only solution is (1, 10).

pooja195 (pooja195):

−2x+y=8 for y: (Add 2x to both sides) \[\huge~\rm~−2x+y+2x=8+2x\] \[\huge~\rm~y=2x+8 \] Substitute 2x+8 for y in 4x+16=2y \[\huge~\rm~4x+16=2(2x+8) \] solve for x

OpenStudy (rebecca1588):

I got 4x+16?

OpenStudy (rebecca1588):

C?

pooja195 (pooja195):

C is correct